OpenStudy (anonymous):

what is the factored form of this trinomial? y^2-11y+30=

OpenStudy (anonymous):

(x - 5)(x-6)

OpenStudy (jamesj):

Trinominal! Fancy work for something so simple. And @Jimmy, perhaps you can explain how you got that.

OpenStudy (anonymous):

yes please explain cuz i dont understan

OpenStudy (anonymous):

ok - i should have explained

OpenStudy (anonymous):

these trinomials which can be factored give two parenthesis with two terms in each. there will be x in each bracket (to make the x^2) and the two numbers are found as follows in this case you need two numbers which when multiplied together give +30 and when added/ subtracted make -11. these are -5 and -6 we can check if we are right by expanding (x-5)(x-6) = x(x-6) - 5(x-6) = x^2 - 6x - 5x + 30 = x^2 -11x + 30

OpenStudy (anonymous):

are you ok with that?

OpenStudy (anonymous):

they are a bit tricky and need some practice before you are confident with them